# Disentangled Interpretable Representation for Efficient Long-term Time Series Forecasting

### v1

@@ III-D Independent Frequential Mapping
III-E Low-rank Weight Sharing
Weight sharing is a common practice in time series forecasting models [27, 29, 19], based on the assumption that different channels of a time series exhibit similar patterns. However, this assumption often does not hold. For example, in Weather datasets, variables such as temperature, pressure, and rainfall may follow distinct patterns. In such cases, training weights independently for each channel can improve performance [28], but this approach significantly increases the number of parameters by a factor of , making the model complexity significantly higher.
To address this, we propose a more efficient alternative. By clustering variables based on their pattern similarities, we can reduce the number of weight sets, balancing the trade-off between model efficiency and performance. This clustering approach not only enhances computational efficiency by reducing the number of learned parameters but also improves accuracy, as the shared weights capture more robust patterns. Specifically, well-structured clustering ensures that the learned representations generalize better across similar variables. Inspired by mixture-of-experts[44] and dynamic convolution[45], we introduce a novel Low-rank Weight Sharing architecture, as shown in Fig. 4, leveraging these clustered patterns to optimize both parameter efficiency and predictive performance.
Specifically, for a multivariate time series with variables, we introduce a hyperparameter , where denotes the number of independent weight sets to be learned, with . For each component, distinct sets of weights are learned, alongside a static routing matrix. The low-rank routing matrix is designed to linearly combine the weights and assign each variable to an appropriate weight set. We regularize the static routing matrix using the Softmax function, where is the Softmax temperature:
@@ III-E Low-rank Weight Sharing
III-F SFALoss
Building upon the insights of FreDF [41], we introduce SFALoss (denoted as ), which combines Weighted Mean Absolute Error (WMAE) in the frequency domain () and MSE in the time domain (). For the frequency domain component (), our SFA module selectively subtracts key frequencies while attenuating noisy frequencies. To refine this process, we modulate the frequency-domain loss through an element-wise multiplication with the SFA weighting factor, , effectively filtering out loss contributions from less relevant frequencies. The WMAE loss in the frequency domain is formulated as follows:
where and are respectively ground truth and predicted future time series in frequency-domain. While computing the loss, is detached from the computational graph, ensuring that no gradients are backpropagated through it.
@@ III-F SFALoss
IV Experiment
IV-A Experimental Setup
IV-A1 Dataset
To comprehensively evaluate the performance and generalization capability of our model, we selected a diverse set of eight datasets, including six for LTSF and two for short-term time series forecasting (STSF). These datasets come from various domains, with different sizes and sampling frequencies, offering a robust benchmark for assessing the versatility and scalability of our approach across a wide range of forecasting tasks. Details are shown in Table I.
For the LTSF task, we utilized six public datasets:
The ETT111https://github.com/zhouhaoyi/ETDataset dataset, including ETTh1, ETTh2, ETTm1 and ETTm2, comprises load and oil temperature data from two power transformers at different stations.
The Electricity222https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 dataset includes electricity consumption data, measured in kilowatt-hours (kWh), from 321 clients, spanning from 2012 to 2014.
The Weather333https://www.bgc-jena.mpg.de/wetter/ dataset consists of 21 meteorological indicators, sampled every 10 minutes, covering the entire year of 2020.
Moreover, to validate the effectiveness of the model in real-world production environments, we selected two workload dataset collected from ByteDance’s cloud servers, serveing as private LTSF datasets:
The FaaS includes 15-day query-per-second (QPS) data from the public cloud platform.
The IaaS includes 45-day CPU usage data from the private cloud platform.
To further demonstrate the generalizability of our model, we extended our evaluation to two short-term forecasting tasks using the following datasets:
The Illness444https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html dataset records weekly influenza-like illness cases as reported by the Centers for Disease Control and Prevention (CDC) in the United States, from 2002 to 2021.
The M5555https://www.kaggle.com/competitions/m5-forecasting-accuracy/ dataset contains hierarchical sales data from Walmart, covering stores in three U.S. states.
| Category | Dataset | Domain | Variates | Samples | Interval |
| LTSF | ETTh1 | Industry | 7 | 17420 | 1 hour |
| ETTh2 | Industry | 7 | 17420 | 1 hour | |
| ETTm1 | Industry | 7 | 69680 | 15 min | |
| ETTm2 | Industry | 7 | 69680 | 15 min | |
| Electricity | Energy | 321 | 26304 | 1 hour | |
| Weather | Environment | 21 | 52696 | 1 hour | |
| FaaS | Cloud Computing | 1014 | 23041 | 1 min | |
| IaaS | Cloud Computing | 168 | 69045 | 1 min | |
| STSF | Illness | Medical | 7 | 966 | 1 week |
| M5 | Market | 30 | 1941 | 1 day |
| DiPE-Linear | FITS | RLinear | DLinear | Nlinear | |||||||
| (Ours) | (ICLR-24) | (arXiv-23) | (AAAI-23) | (AAAI-23) | |||||||
@@ IV-A1 Dataset
| DiPE-Linear | iTransformer | PatchTST† | TimesNet† | Crossformer† | MICN† | FEDformer† | |||||||||
| (Ours) | (ICLR-24) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICML-22) | |||||||||
| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| ETTh1 | 96 | 0.369 | 0.393 | 0.386 | 0.405 | 0.370 | 0.399 | 0.384 | 0.402 | 0.386 | 0.429 | 0.396 | 0.427 | 0.376 | 0.415 |
| 192 | 0.407 | 0.415 | 0.424 | 0.440 | 0.413 | 0.421 | 0.557 | 0.436 | 0.419 | 0.444 | 0.430 | 0.453 | 0.423 | 0.446 | |
| 336 | 0.424 | 0.427 | 0.449 | 0.460 | 0.422 | 0.436 | 0.491 | 0.469 | 0.440 | 0.461 | 0.433 | 0.458 | 0.444 | 0.462 | |
| 720 | 0.409 | 0.439 | 0.495 | 0.487 | 0.447 | 0.466 | 0.521 | 0.500 | 0.519 | 0.524 | 0.474 | 0.508 | 0.469 | 0.492 | |
| ETTh2 | 96 | 0.275 | 0.336 | 0.297 | 0.348 | 0.274 | 0.336 | 0.340 | 0.374 | 0.276 | 0.338 | 0.289 | 0.357 | 0.332 | 0.374 |
| 192 | 0.325 | 0.372 | 0.371 | 0.403 | 0.339 | 0.379 | 0.402 | 0.414 | 0.345 | 0.382 | 0.409 | 0.438 | 0.407 | 0.446 | |
| 336 | 0.350 | 0.393 | 0.404 | 0.428 | 0.329 | 0.380 | 0.452 | 0.452 | 0.375 | 0.411 | 0.417 | 0.452 | 0.400 | 0.447 | |
| 720 | 0.375 | 0.415 | 0.424 | 0.444 | 0.379 | 0.422 | 0.462 | 0.468 | 0.408 | 0.446 | 0.426 | 0.473 | 0.412 | 0.469 | |
| ETTm1 | 96 | 0.309 | 0.350 | 0.300 | 0.353 | 0.290 | 0.342 | 0.338 | 0.375 | 0.316 | 0.373 | 0.314 | 0.360 | 0.326 | 0.390 |
| 192 | 0.339 | 0.369 | 0.345 | 0.382 | 0.332 | 0.369 | 0.371 | 0.387 | 0.377 | 0.411 | 0.359 | 0.387 | 0.365 | 0.415 | |
| 336 | 0.367 | 0.386 | 0.374 | 0.398 | 0.369 | 0.392 | 0.410 | 0.411 | 0.431 | 0.442 | 0.398 | 0.413 | 0.392 | 0.425 | |
| 720 | 0.416 | 0.413 | 0.429 | 0.430 | 0.416 | 0.420 | 0.478 | 0.450 | 0.600 | 0.547 | 0.459 | 0.464 | 0.446 | 0.458 | |
| ETTm2 | 96 | 0.162 | 0.252 | 0.175 | 0.266 | 0.165 | 0.255 | 0.187 | 0.267 | 0.421 | 0.461 | 0.178 | 0.273 | 0.180 | 0.271 |
| 192 | 0.216 | 0.289 | 0.242 | 0.312 | 0.220 | 0.292 | 0.249 | 0.309 | 0.503 | 0.519 | 0.245 | 0.316 | 0.252 | 0.318 | |
| 336 | 0.268 | 0.324 | 0.282 | 0.340 | 0.274 | 0.329 | 0.321 | 0.351 | 0.611 | 0.580 | 0.295 | 0.350 | 0.324 | 0.364 | |
| 720 | 0.353 | 0.379 | 0.378 | 0.398 | 0.362 | 0.385 | 0.497 | 0.403 | 0.996 | 0.750 | 0.389 | 0.406 | 0.410 | 0.420 | |
| ECL | 96 | 0.132 | 0.228 | 0.132 | 0.229 | 0.129 | 0.222 | 0.168 | 0.272 | 0.187 | 0.283 | 0.159 | 0.267 | 0.186 | 0.302 |
| 192 | 0.148 | 0.245 | 0.151 | 0.246 | 0.147 | 0.240 | 0.184 | 0.289 | 0.258 | 0.330 | 0.168 | 0.279 | 0.197 | 0.311 | |
| 336 | 0.162 | 0.261 | 0.167 | 0.264 | 0.163 | 0.259 | 0.198 | 0.300 | 0.323 | 0.369 | 0.196 | 0.308 | 0.213 | 0.328 | |
| 720 | 0.198 | 0.296 | 0.194 | 0.286 | 0.197 | 0.290 | 0.220 | 0.320 | 0.404 | 0.423 | 0.203 | 0.312 | 0.233 | 0.344 | |
| Weather | 96 | 0.142 | 0.201 | 0.159 | 0.208 | 0.149 | 0.198 | 0.172 | 0.220 | 0.153 | 0.217 | 0.161 | 0.226 | 0.238 | 0.314 |
| 192 | 0.187 | 0.253 | 0.200 | 0.248 | 0.194 | 0.241 | 0.219 | 0.261 | 0.197 | 0.269 | 0.220 | 0.283 | 0.275 | 0.329 | |
| 336 | 0.234 | 0.293 | 0.253 | 0.289 | 0.245 | 0.282 | 0.280 | 0.306 | 0.252 | 0.311 | 0.275 | 0.328 | 0.339 | 0.377 | |
| 720 | 0.306 | 0.348 | 0.321 | 0.338 | 0.314 | 0.334 | 0.365 | 0.359 | 0.318 | 0.363 | 0.311 | 0.356 | 0.389 | 0.409 | |
| Best count | 16 | 14 | 1 | 1 | 8 | 11 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
IV-A2 Implementation details
All experiments are implemented in PyTorch [46] on an NVIDIA 4090D GPU. The models are trained using the Adam optimizer [47] with a learning rate of for epochs. Batch sizes are set to 64 and 8 for LTSF and STSF datasets, respectively. For Electricity and Weather datasets, , and otherwise. Input sequence lengths are , , and for public LTSF, private LTSF, and STSF datasets, respectively. Following [45], we apply linear annealing for the softmax temperature in Low-rank Weight Sharing to ensure balanced learning during early epochs.
IV-A3 Baselines
In this work, we focus on evaluating and comparing linear models, benchmarking our approach against leading methods such as FITS [29], RLinear [28], DLinear, and NLinear [27]. To further demonstrate the effectiveness of our model, we also compare it with nonlinear deep learning models, including iTransformer [20], PatchTST [48], TimesNet [11], Crossformer [18], MICN [10], and FEDformer [17], showing that our model achieves state-of-the-art or comparable performance across both linear and nonlinear baselines. Additionally, SparseTSF [37] and FITS [29] are selected as parameter-efficient baselines to highlight the efficiency of our model design.
IV-A4 Evaluation Metrics
We follow the established practices in prior works to evaluate forecasting performance using two primary metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE).
IV-B Main Results
IV-B1 Versus FCs
We first compared our model DiPE-Linear with other leading FCs. To ensure a fair evaluation, all models were trained using the same trainer settings. Throughout the training process, we carefully monitored each model to ensure convergence under these consistent conditions. All models use the optimal model hyperparameters provided in the authors’ original implementations. Table II presents the evaluation results for the linear models on the LTSF and STSF datasets.
@@ IV-B1 Versus FCs
One notable advantage of DiPE-Linear is its substantially reduced parameter count compared to FCs. By operating within a low-dimensional manifold in the FCs’ model space, DiPE-Linear achieves performance comparable to or better than FCs, suggesting that our carefully designed model space contains the optimal solution to TSF tasks. This highlights the robustness and effectiveness of our architecture in representing the essential solution space of FCs with minimal complexity.
IV-B2 Versus nonlinear models
@@
Next, we compare DiPE-Linear with several nonlinear models, including Transformer-based models amd CNN-based models. Table III presents the evaluation results for these nonlinear models on public LTSF datasets. For private LTSF datasets, excessively long input lengths result in high memory usage for nonlinear models, leading to unacceptably high deployment costs in production environments. For STSF datasets, the limited amount of training data makes the models highly prone to overfitting, making training challenging. Therefore, we do not report model performance on these two types of datasets. To ensure a fair comparison, all nonlinear models were trained with input lengths , and the best results were selected for evaluation.
Our model demonstrates a slight performance advantage over PatchTST, while requiring only times the parameter count. Furthermore, our model significantly outperforms other nonlinear models, highlighting its competitive advantages in both efficiency and performance.
IV-B3 Parameter complexity and efficiency
Recent studies have focused on parameter-efficient models, and we selected SparseTSF and FITS as baselines. In the FITS model, which employs a low-pass filter, we set the cutoff frequencies to the maximum and minimum harmonics provided in its source code, denoting them as FITS-T and FITS-L, respectively. Experiments on ETTh1, ETTm1, Electricity, and Weather datasets (Fig. 5) show that our model achieves the lowest MSE overall, with a parameter count falling between FITS-T and SparseTSF. While FITS-T reduces parameters by blocking more high-frequency information, it sacrifices performance. In contrast, our model not only uses fewer parameters than FITS-T but also matches or outperforms FITS-L. This demonstrates that our model design effectively balances performance and parameter efficiency for time series forecasting.
Table IV summarizes the parameter counts of leading models on the LTSF dataset, using an input look-back length of and an output length of . DiPE-Linear emerges as the most lightweight model, with parameter counts reduced by a factor of to compared to other baseline models across four LTSF datasets. This compact design not only significantly lowers computational requirements but also aligns with environmentally sustainable practices by reducing the carbon footprint.
| Model | ETTh1 | ETTm1 | Electricity | Weather |
| DiPE-Linear (Ours) | 4K | 4K | 17K | 16K |
| FITS | 154K | 60K | 411K | 572K |
| NLinear | 519K | 519K | 519K | 167M |
| RLinear | 519K | 519K | 519K | 167M |
| DLinear | 1M | 1M | 1M | 333M |
| iTransformer | 1M | 385K | 5M | 5M |
| TimesNet | 1M | 1M | 151M | 2M |
| PatchTST | 36M | 42M | 40M | 30M |
| Crossformer | 43M | 43M | 18M | 216K |
@@ IV-B3 Parameter complexity and efficiency
IV-B4 Effect of Training Data Size
To further investigate whether the model can perform effectively with reduced training data, we conducted experiments on the ETTh1 and ETTm1 datasets using different proportions of the training set. For the ETTh1 dataset, we compared our model with other FCs. Additionally, since the ETTm1 dataset contains more samples, we included nonlinear models for a broader comparison. As shown in Fig. 6, DiPE-Linear outperforms both FCs and nonlinear models across all proportions of the training set. Notably, on the ETTh1 dataset, our model surpasses the performance of other FC models trained on the full dataset while using only 40% of the training data.
IV-C Ablation Studies
IV-C1 Basic Components
To comprehensively validate the effectiveness of our model, we conducted ablation tests on two submodules: SFA and STA. The submodule IFM, which maps the input sequence space to the output sequence space, serves as the core component of our model and therefore cannot be ablated. Table V presents the results of the ablation experiments on the ETTh2 and Weather datasets with various prediction lengths. In both the ETTh2 and Weather datasets, the incorporation of the SFA and STA modules generally enhances the model’s prediction accuracy. Specifically, in the Weather dataset, the abundance of training samples ensures that each submodule effectively captures significant patterns, leading to a substantial improvement in prediction performance. This demonstrates the necessity of addressing the importance of both the frequency domain and time domain in our input data processing.

### v2

@@ 3.4 Independent Frequential Mapping
3.5 Low-rank Weight Sharing
Weight sharing is commonly used in time series forecasting models [23, 21, 14] under the assumption that different channels exhibit similar patterns. However, this assumption does not hold in datasets like Weather, where variables such as temperature, pressure, and rainfall follow distinct patterns. While training weights independently for each channel can improve performance [7], it significantly increases model complexity. To address this, we propose selecting a lower rank based on pattern similarities, reducing the number of weight sets and balancing efficiency with performance. Inspired by mixture-of-experts [13] and dynamic convolution [2], we introduce a novel Low-rank Weight Sharing architecture, optimizing parameter efficiency and predictive performance.
Specifically, for a multivariate time series with variables, we introduce a hyperparameter , where denotes the number of independent weight sets to be learned, with . For each component, distinct sets of weights are learned, alongside a learnable routing matrix. The low-rank routing matrix is a model parameter designed to linearly combine the weights and assign each variable to an appropriate weight set. We regularize this routing matrix using the Softmax function, where is the Softmax temperature:
@@ 3.5 Low-rank Weight Sharing
3.6 SFALoss
Building on the insights from FreDF [17], we introduce SFALoss (denoted as ), which integrates the Weighted Mean Absolute Error (WMAE) in the frequency domain () and Mean Squared Error (MSE) in the time domain (). We empirically found this combination to be most effective: WMAE for provides a more robust optimization signal that is less sensitive to large magnitude errors in a few dominant frequencies, while the standard MSE for ensures overall predictive accuracy in the time domain. The frequency-domain loss () is modulated by element-wise multiplication with the SFA weighting factor, , allowing prioritizing relevant frequencies:
where and are respectively ground truth and predicted future time series in frequency-domain. While computing the loss, is detached from the computational graph, ensuring that no gradients are backpropagated through it. This crucial step prevents a trivial solution where the SFA module might learn to suppress (i.e., set weights to zero for) frequencies that are difficult to predict, rather than learning the frequencies most relevant to the primary time-domain forecasting task.
@@ 3.6 SFALoss
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets.
We evaluate our model on 10 real-world datasets, comprising 8 for Long-term (LTSF) and 2 for Short-term (STSF) forecasting. The LTSF datasets include ETT (ETTh1, ETTh2, ETTm1, ETTm2) [24], Weather [24], Electricity, and two private cloud service datasets (IaaS, FaaS). The STSF datasets are M5 and Illness [20]. Further preprocessing details are available in the accompanying code.
4.1.2 Implementation Details.
Models are trained for 50 epochs with a learning rate. Batch sizes are set to 64 (LTSF) and 8 (STSF). Forecasting horizons are for LTSF and for STSF. Input sequence lengths are 720 (public LTSF), 2880 (private LTSF), and 60 (STSF). We set for Electricity and Weather, otherwise, and apply linear annealing for the softmax temperature [2].
4.1.3 Baselines and Metrics.
We benchmark our approach against leading linear models, including FITS [21], RLinear [7], DLinear, and NLinear [23]. We also compare with established non-linear models (e.g., iTransformer [11], PatchTST [3], TimesNet [19]) and lightweight baselines (SparseTSF [10], FITS [21]). Performance is evaluated using MSE and MAE.
4.2 Main Results
4.2.1 Versus FC Models.
| DiPE-Linear (Ours) | FITS (2024) | RLinear (2023) | DLinear (2023) | Nlinear (2023) | |||||||
@@ 4.2 Main Results
We first compared our model DiPE-Linear with other leading FC models. To ensure a fair evaluation, all models were trained to converge under consistent training settings and using optimal hyperparameters from the authors’ original implementations. Table 1 presents the evaluation results for FC models on the LTSF and STSF datasets.
@@
One notable advantage of DiPE-Linear is its substantially reduced parameter count compared to FC models. By operating within a low-dimensional manifold in the FC models’ model space, DiPE-Linear achieves performance comparable to or better than FC models, suggesting that our carefully designed model space provides a highly efficient and effective solution for these LTSF benchmark tasks. This highlights the robustness and effectiveness of our architecture in representing the essential solution space of FC models with minimal complexity.
4.2.2 Versus nonlinear models.
| DiPE-Linear | TimeMixer++ | TQNet | iTransformer | PatchTST | TimesNet | Crossformer | MICN | FEDformer | |||||||||||
| (Ours) | (2025) | (2025) | (2024) | (2023) | (2023) | (2023) | (2023) | (2022) | |||||||||||
| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| ETTh1 | 96 | 0.369 | 0.393 | 0.361 | 0.403 | 0.371 | 0.393 | 0.386 | 0.405 | 0.370 | 0.399 | 0.384 | 0.402 | 0.386 | 0.429 | 0.396 | 0.427 | 0.376 | 0.415 |
| 192 | 0.407 | 0.415 | 0.416 | 0.441 | 0.428 | 0.426 | 0.424 | 0.440 | 0.413 | 0.421 | 0.557 | 0.436 | 0.419 | 0.444 | 0.430 | 0.453 | 0.423 | 0.446 | |
| 336 | 0.424 | 0.427 | 0.430 | 0.434 | 0.476 | 0.446 | 0.449 | 0.460 | 0.422 | 0.436 | 0.491 | 0.469 | 0.440 | 0.461 | 0.433 | 0.458 | 0.444 | 0.462 | |
| 720 | 0.409 | 0.439 | 0.467 | 0.451 | 0.487 | 0.470 | 0.495 | 0.487 | 0.447 | 0.466 | 0.521 | 0.500 | 0.519 | 0.524 | 0.474 | 0.508 | 0.469 | 0.492 | |
| ETTh2 | 96 | 0.275 | 0.336 | 0.276 | 0.328 | 0.295 | 0.343 | 0.297 | 0.348 | 0.274 | 0.336 | 0.340 | 0.374 | 0.276 | 0.338 | 0.289 | 0.357 | 0.332 | 0.374 |
| 192 | 0.325 | 0.372 | 0.342 | 0.379 | 0.367 | 0.393 | 0.371 | 0.403 | 0.339 | 0.379 | 0.402 | 0.414 | 0.345 | 0.382 | 0.409 | 0.438 | 0.407 | 0.446 | |
| 336 | 0.350 | 0.393 | 0.346 | 0.398 | 0.417 | 0.427 | 0.404 | 0.428 | 0.329 | 0.380 | 0.452 | 0.452 | 0.375 | 0.411 | 0.417 | 0.452 | 0.400 | 0.447 | |
| 720 | 0.375 | 0.415 | 0.392 | 0.415 | 0.433 | 0.446 | 0.424 | 0.444 | 0.379 | 0.422 | 0.462 | 0.468 | 0.408 | 0.446 | 0.426 | 0.473 | 0.412 | 0.469 | |
| ETTm1 | 96 | 0.309 | 0.350 | 0.310 | 0.334 | 0.311 | 0.353 | 0.300 | 0.353 | 0.290 | 0.342 | 0.338 | 0.375 | 0.316 | 0.373 | 0.314 | 0.360 | 0.326 | 0.390 |
| 192 | 0.339 | 0.369 | 0.348 | 0.362 | 0.356 | 0.378 | 0.345 | 0.382 | 0.332 | 0.369 | 0.371 | 0.387 | 0.377 | 0.411 | 0.359 | 0.387 | 0.365 | 0.415 | |
| 336 | 0.367 | 0.386 | 0.376 | 0.391 | 0.390 | 0.401 | 0.374 | 0.398 | 0.369 | 0.392 | 0.410 | 0.411 | 0.431 | 0.442 | 0.398 | 0.413 | 0.392 | 0.425 | |
| 720 | 0.416 | 0.413 | 0.440 | 0.423 | 0.452 | 0.440 | 0.429 | 0.430 | 0.416 | 0.420 | 0.478 | 0.450 | 0.600 | 0.547 | 0.459 | 0.464 | 0.446 | 0.458 | |
| ETTm2 | 96 | 0.162 | 0.252 | 0.170 | 0.245 | 0.173 | 0.256 | 0.175 | 0.266 | 0.165 | 0.255 | 0.187 | 0.267 | 0.421 | 0.461 | 0.178 | 0.273 | 0.180 | 0.271 |
| 192 | 0.216 | 0.289 | 0.229 | 0.291 | 0.238 | 0.298 | 0.242 | 0.312 | 0.220 | 0.292 | 0.249 | 0.309 | 0.503 | 0.519 | 0.245 | 0.316 | 0.252 | 0.318 | |
| 336 | 0.268 | 0.324 | 0.303 | 0.343 | 0.301 | 0.340 | 0.282 | 0.340 | 0.274 | 0.329 | 0.321 | 0.351 | 0.611 | 0.580 | 0.295 | 0.350 | 0.324 | 0.364 | |
| 720 | 0.353 | 0.379 | 0.373 | 0.399 | 0.397 | 0.396 | 0.378 | 0.398 | 0.362 | 0.385 | 0.497 | 0.403 | 0.996 | 0.750 | 0.389 | 0.406 | 0.410 | 0.420 | |
| Electricity | 96 | 0.132 | 0.228 | 0.135 | 0.222 | 0.134 | 0.229 | 0.132 | 0.229 | 0.129 | 0.222 | 0.168 | 0.272 | 0.187 | 0.283 | 0.159 | 0.267 | 0.186 | 0.302 |
| 192 | 0.148 | 0.245 | 0.147 | 0.235 | 0.154 | 0.247 | 0.151 | 0.246 | 0.147 | 0.240 | 0.184 | 0.289 | 0.258 | 0.330 | 0.168 | 0.279 | 0.197 | 0.311 | |
| 336 | 0.162 | 0.261 | 0.164 | 0.245 | 0.169 | 0.264 | 0.167 | 0.264 | 0.163 | 0.259 | 0.198 | 0.300 | 0.323 | 0.369 | 0.196 | 0.308 | 0.213 | 0.328 | |
| 720 | 0.198 | 0.296 | 0.212 | 0.310 | 0.201 | 0.294 | 0.194 | 0.286 | 0.197 | 0.290 | 0.220 | 0.320 | 0.404 | 0.423 | 0.203 | 0.312 | 0.233 | 0.344 | |
| Weather | 96 | 0.142 | 0.201 | 0.155 | 0.205 | 0.157 | 0.200 | 0.159 | 0.208 | 0.149 | 0.198 | 0.172 | 0.220 | 0.153 | 0.217 | 0.161 | 0.226 | 0.238 | 0.314 |
| 192 | 0.187 | 0.253 | 0.201 | 0.245 | 0.206 | 0.245 | 0.200 | 0.248 | 0.194 | 0.241 | 0.219 | 0.261 | 0.197 | 0.269 | 0.220 | 0.283 | 0.275 | 0.329 | |
| 336 | 0.234 | 0.293 | 0.237 | 0.265 | 0.262 | 0.287 | 0.253 | 0.289 | 0.245 | 0.282 | 0.280 | 0.306 | 0.252 | 0.311 | 0.275 | 0.328 | 0.339 | 0.377 | |
| 720 | 0.306 | 0.348 | 0.312 | 0.334 | 0.344 | 0.342 | 0.321 | 0.338 | 0.314 | 0.334 | 0.365 | 0.359 | 0.318 | 0.363 | 0.311 | 0.356 | 0.389 | 0.409 | |
| Best count | 15 | 11 | 2 | 9 | 0 | 1 | 1 | 1 | 8 | 5 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
@@
Next, we compare DiPE-Linear with several nonlinear models. Table 2 presents the evaluation results for these nonlinear models specifically on the public LTSF datasets. We do not report model performance on the private LTSF and STSF datasets due to the following practical limitations: for private LTSF datasets, excessively long input lengths lead to high memory usage for nonlinear models, resulting in prohibitively high deployment costs in production environments. For STSF datasets, the limited amount of training data makes these models highly prone to overfitting, which complicates the training process.
To ensure a rigorous comparison, baseline models were evaluated by selecting the best result from a comprehensive search over input lengths . To benchmark against the very latest advancements, we also include two recent models, TimeMixer++[18] and TQNet[9], citing the best performance reported in their original publications.
The results in Table 2 highlight the remarkable effectiveness of DiPE-Linear. Notably, our model establishes a new state-of-the-art, achieving the best overall accuracy against all competing methods. It significantly outperforms established baselines like iTransformer and recent models such as TimeMixer++[18] and TQNet[9]. This SOTA performance is achieved with unparalleled efficiency. For instance, DiPE-Linear matches or exceeds the accuracy of PatchTST [3] while utilizing only of its parameters. This massive reduction in model complexity, combined with its superior forecasting accuracy, underscores the competitive advantage of DiPE-Linear and establishes it as a new, highly efficient benchmark for LTSF.
Furthermore, our state-of-the-art results, achieved with a purely linear model, contribute strong evidence to the ongoing debate [23] regarding the necessity of complex, non-linear architectures for LTSF. The success of DiPE-Linear suggests that, for these benchmarks, a well-structured inductive bias (i.e., the factorization of frequential and temporal components) is more critical for generalization than the capacity for high-dimensional non-linear mapping (as found in Transformers).
4.2.3 Parameter complexity and efficiency.
Recent studies have focused on parameter efficient models. We compare DiPE-Linear with SparseTSF [10] and FITS [21] (FITS-T and FITS-L, representing different low-pass filter settings). As shown in Fig. 4, our model achieves the lowest MSE across these parameter-efficient baselines. Notably, DiPE-Linear not only uses fewer parameters than FITS-T but also matches or outperforms FITS-L, demonstrating a superior balance of performance and parameter efficiency.
@@
4.2.4 Effect of Training Data Size.
To further investigate whether the model can perform effectively with reduced training data, we conducted experiments on the ETTh1 and ETTm1 datasets using different proportions of the training set. For the ETTh1 dataset, we compared our model with other FC models. Additionally, since the ETTm1 dataset contains more samples, we included nonlinear models for a broader comparison. As shown in Fig. 5, DiPE-Linear outperforms both FC models and nonlinear models across all proportions of the training set. Notably, on the ETTh1 dataset, our model surpasses the performance of other FC models trained on the full dataset while using only 40% of the training data.
4.3 Ablation Studies
4.3.1 Basic Components.
To comprehensively validate the effectiveness of our model, we conducted ablation tests on two submodules: SFA and STA. The submodule IFM, which maps the input sequence space to the output sequence space, serves as the core component of our model and therefore cannot be ablated. Table 3 presents the results of the ablation experiments on the ETTh2 and Weather datasets with various prediction lengths. In both the ETTh2 and Weather datasets, the incorporation of the SFA and STA modules generally enhances the model’s prediction accuracy. Specifically, in the Weather dataset, the abundance of training samples ensures that each submodule effectively captures significant patterns, leading to a substantial improvement in prediction performance. This demonstrates the necessity of addressing the importance of both the frequency domain and time domain in our input data processing.